OpenStudy (anonymous):

x + 2y - z = -1 x - 4y - z = -4 2x - 2y + 3z = 15

OpenStudy (anonymous):

solve using matrices

OpenStudy (anonymous):

http://www.mathportal.org/calculators/system-of-equations-solver/system-3x3.php Heres a website that will solve your 3 system equation let me know if you can access it to double-check your answer.

OpenStudy (anonymous):

I dont understand the explanation it gave

OpenStudy (anonymous):

Sorry @cookiimonster27 did you get it? Sorry my typing came out weird... Did you put in your equation, click, and then solve using Cramers rule??

OpenStudy (anonymous):

I checked it but it doesnt make any sense let me try again

OpenStudy (anonymous):

Still doesn't make sense?

OpenStudy (anonymous):

@cookiimonster27 do you still need help or do you get it?

OpenStudy (mathmath333):

\(\large\tt \color{black}{ \left( \begin{array}{ccc} 1 & 2 & -1 \\ 1 & -4 & -1 \\ 2 & -2 & 3 \end{array} \right) \times \left( \begin{array}{ccc} x \\ y \\ z \end{array} \right)=\left( \begin{array}{ccc} -1 \\ -4 \\ 15 \end{array} \right)}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{ [A][X]=[B]}\) \(\large\tt \color{black}{ [X]=[A^{-1}][B]}\)

OpenStudy (mathmath333):

U have to find inverse of \(\large\tt \color{black}{A=\left( \begin{array}{ccc} 1 & 2 & -1 \\ 1 & -4 & -1 \\ 2 & -2 & 3 \end{array} \right),A^{-1}=?}\)

OpenStudy (mathmath333):

and multiply it into \(\large\tt \color{black}{[B]=\left( \begin{array}{ccc} -1 \\ -4 \\ 15 \end{array} \right)}\)

OpenStudy (mathmath333):

then u will get \(\large\tt \color{black}{[X]=\left( \begin{array}{ccc} ? \\ ? \\ ? \end{array} \right)}\)